Hahn Manufacturing purchases a key component of one of its products from a local supplier. The current purchase price is $ 1 comma 450 per unit. Efforts to standardize parts succeeded to the point that this same component can now be used in five different products. Annual component usage should increase from 150 to 700 units. Management wonders whether it is time to make the component​ in-house, rather than to continue buying it from the supplier. Fixed costs would increase by about $ 40 comma 000 per year for the new equipment and tooling needed. The cost of raw materials and variable overhead would be about $ 1 comma 000 per​ unit, and labor costs would be $ 350 per unit produced. a. Should Hahn make rather than​ buy? Hahn should ▼ buy make the​ components, saving ​$ nothing per year as compared to the other decision. ​(Enter your response rounded to the nearest whole​ number.)
Answer #1

Cost that can be incurred from local supplier = number of units procured * cost per unit

= 700 * 1450 = $1,015,000

Cost that can be incurred if it is made in house:

Total cost = Total fixed cost + total variable cost

Total fixed cost = $40,000

Total variables = 1,000 + 350 = $1,350

Total variable cost = 1350 * 700 = $945,000

Total cost = $40,000 + $945,000 = $985,000

As the cost of in house making is less than the cost of local supplier, Hahn can go for inhouse making to save $30,000.